Here is a word problem. Write a system of two equations in two unknowns and solve by the method of your choice. Don’t forget to show the methodology. Irving has 35 coins, consisting of only dimes and nickels. If the value of his coins is $3.30, how many nickels and dimes does he have?

Answer by mananth(16949)

let the number of dimes be x
& the number of nickels be y
x+y=35
value of the coins
10x+5y=330 cents
10(x+y)-(10x+5y) = 10*35 -330
10x+10y-10x-5y=350-330
5y=20
y=4 the number of nickels
35-4 =31 the number of dimes
ANANTH